# Space spanned by characteristic exponents

### v2

@@ 1.6.1. The Zariski-dense orbit conjecture
By applying Corollary 1.7 we can give a new proof of a special case of the Zariski-dense orbit conjecture.
@@
Let $k$ be an algebraically closed field of characteristic 00. Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$. If we have $\{g\in k(X):g\circ f=g\}=k$ where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
@@
The converse of ZDO holds and is easy. For some progress on ZDO, see [1, 2, 17, 27, 29] for example.
@@
Let $X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}$ be the variety of product of $N$ copies of projective line over an algebraically closed field $k$ of characteristic 00. Suppose that $f:X\to X$ is an endomorphism of the form $f_{1}\times\cdots\times f_{N}$ where $f_{j}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ is a non-constant rational map for $1\leq j\leq N$. The ZDO holds for $(X,f)$.
@@
The original proof of Theorem 1.12 in [29] relies on the solution of the (adelic) Zariski-dense orbit conjecture on smooth projective surfaces [29, Theorem 1.15], the notion of adelic topology introduced in [29, §3] and a classification result [29, Proposition 9.2] on invariant subvarieties of $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ (see also [17, 8]). When $n=2$, Pakovich gave another proof [22] of Theorem 1.12 using his classification of invariant curves in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and some height argument. We give a new proof in the case where all the $f_{i}$’s are non-exceptional of degree at least $2$, and this proof does not need the ingredients mentioned above. Other cases of Theorem 1.12 are relatively easy or can be reduced to the case where $f_{i}$ are all non-exceptional, as shown in [29, §9.3].
@@ 1.6.2. A characterization of PCF maps
Here $K_{x}$ (resp. $L_{x}$) is any number field containing $\rho_{f}(x)$ (resp. $\lvert\rho_{f}(x)\rvert$) and $N_{K_{x}/\mathbb{Q}}$ (resp. $N_{L_{x}/\mathbb{Q}}$) is the norm map for the extension $K_{x}/\mathbb{Q}$ (resp. $L_{x}/\mathbb{Q}$), i.e., the determinant of the $\mathbb{Q}$-linear transformation induced by multiplication by $\rho_{f}(x)$ (resp. $|\rho_{f}(x)|$).
Clearly, the subspaces $V,W$ above are independent of the choices of the fields $K_{x},L_{x}$, respectively. We also remark that the $\mathbb{Q}$-vector spaces $V$ in $(2)$ and $W$ in $(3)$ are the same space.
@@
Similarly, the actual proof shows that we can weaken the conditions $(2)$ and $(3)$ in Theorem 1.14, in the spirit of Remarks 1.3 and 1.6. Precisely, we can add two more equivalent statements in Theorem 1.14:
$(2)^{*}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V^{*}=V^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $V^{*}$ is generated over $\mathbb{Q}$ by
@@
$(3)^{*}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all but finitely $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W^{*}=W^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $W^{*}$ is generated over $\mathbb{Q}$ by
@@ 2. Rational maps with algebraic lengths
Let $K$ be an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation $\tau$, i.e., $\tau(K)=K$. The aim of this section is to prove the following theorem.
@@
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $\lvert\rho_{f}(x)\rvert\in K$. Then $f$ is defined over $K$.
@@
In fact, we can weaken the condition of Theorem 2.1 by only requiring that $f$ is not flexible Lattès and $\lvert\rho_{f}(x)\rvert\in K$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$. To show this, we need to replace $A$ in the proof in §2.7 by $A=(A_{n}=(RL^{*}(f)_{n})\cap K))_{n}$ and then the original proof works. Here, $A_{n}=(RL^{*}(f)_{n})\cap K$ is the sub-multiset of the multiset $RL^{*}(f)_{n}$ of all elements also in $K$, see §2.6 and §2.7.
Applying Theorem 2.1 to the case $K=\overline{\mathbb{Q}}$, we get Theorem 1.2. We will introduce the notions of Weil restriction, admissible subsets, and transcendental points in §2.1-2.3. Then we construct certain moduli spaces and study the length map on them in §2.4-2.6. After these preparations, we finish the proof of Theorem 2.1 in §2.7.
@@ 2.1. Weil restriction
and both $\frac{u+\tau(u)}{2}$ and $\frac{u-\tau(u)}{2i}$ are contained in $L$. This concludes the proof. ∎
In order to prove Theorem 2.1, we need to analyze the length map respecting the structure of $\mathcal{M}_{d}(K)$ as an algebraic variety over $L$ using Weil restriction, due to the appearance of the absolute value in the condition. Hence, the notion of Weil restriction, which we will briefly recall, is necessary. See [23, §4.6] and [4, §7.6] for more information.
Denote by ${\rm Var}_{/K}$ (resp. ${\rm Var}_{/L}$) the category of varieties over $K$ (resp. $L$). For every variety $X$ over $K$, there is a unique variety $R(X)$ over $L$ which represents the functor ${\rm Var}_{/L}\to{\rm Sets}$ sending $V\in{\rm Var}_{/L}$ to ${\rm Hom}(V\otimes_{L}K,X)$. It is called the Weil restriction of $X$ with respect to $K/L$. The functor $X\mapsto R(X)$ is called the Weil restriction. One has the canonical bijection $\psi_{K}:X(K)\to R(X)(L)$ between sets. When $K=\mathbb{C}$, this map is a real analytic diffeomorphism. One may view $X(K)$ as the set of $L$-points of an $L$-algebraic variety $R(X)$ via $\psi_{K}$.
@@
The $L$-Zariski topology on $X(K)$ is the restriction of the Zariski topology on $R(X)$ to $R(X)(L)$ via the identification $\psi_{K}$. A subset $Y$ of $X(K)$ is called $L$-algebraic if it is closed in the $L$-Zariski topology. When $K=\mathbb{C}$, the $L$-Zariski topology is exactly the real Zariski topology as in [12, §8.1.1].
By the last statement of Proposition 2.6 below, the $L$-Zariski topology is stronger than the Zariski topology on $X(K)$.
@@
Let $X,Y\in{\rm Var}_{/K}$, then we have the following properties:
@@
If $X$ is the subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the equations $\sum_{I}a_{i,I}z^{I}=0$, $i=1,\dots,s$. Then $X^{\tau}$ is the subvariety of $\mathbb{A}^{N}_{K}$ defined by $\sum_{I}\tau(a_{i,I})z^{I}=0$, $i=1,\dots,s$. The map $\tau:X=(X^{\tau})^{\tau}\to X^{\tau}$ sends a point $(z_{1},\dots,z_{N})\in X(K)$ to $(\tau(z_{1}),\dots,\tau(z_{N}))\in X^{\tau}(K)$.
@@
Proposition 2.8.
[23, Exercise 4.7] We have a canonical isomorphism
@@ 2.2. Admissible subsets
In this subsection, we recall the notion of admissible subsets of real algebraic varieties introduced in [12], which satisfies the important decreasing chain condition.
@@
Definition 2.9.
[12, §8.2] A closed subset $V$ of $X(\mathbb{R})$ (with the real topology) is called admissible if there is a morphism $f:Y\to X$ of real algebraic varieties and a Zariski closed subset $V^{\prime}\subseteq Y$ such that $V=f(V^{\prime}(\mathbb{R}))$ and $f$ is étale at every point in $V^{\prime}(\mathbb{R})$.
@@
Denote by $J$ the non-étale locus for $f$ in $Y$. We have $J\cap V^{\prime}(\mathbb{R})=\emptyset$. Since we may replace $(Y,V^{\prime})$ by $(Y\setminus J,V^{\prime}\cap(Y\setminus J))$, in the above definition we may further assume that $f$ is étale.
Proposition 2.11.
[12, Remarks 8.14, 8.15, and Proposition 8.16] We have the following basic properties:
@@
Theorem 2.12.
[12, Theorem 8.17] Let $(V_{n})_{n\geq 1}$ be a sequence of decreasing admissible subsets of $X(\mathbb{R})$. Then there is $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
@@ 2.3. Transcendental points
Let $X_{K}$ be a variety over $K$ and $X:=X_{K}\otimes_{K}\mathbb{C}$. We consider $X_{K}$ as a model of $X$ over $K$.
Denote by $\pi_{K}:X\to X_{K}$ the natural projection. For every point $x\in X(\mathbb{C})$, define $Z(x)_{K}$ to be the Zariski closure of $\pi_{K}(x)$ in $X_{K}$ and $Z(x):=\pi_{K}^{-1}(Z(x)_{K})\subseteq X$. It is clear that $Z(x)$ is irreducible. We call $Z(x)$ the $\mathbb{C}/K$-closure of $x$ w.r.t the model $X_{K}$. We say that $x$ is transcendental if $\dim Z(x)\geq 1$ and call $\dim Z(x)$ the transcendental degree of $x$.
The notion of transcendental points plays important role in the proof of some cases of the geometric Bombieri-Lang conjecture by Xie-Yuan [30, 31] and the proof of the dynamical André-Oort conjecture for curves by Ji-Xie [11]. Roughly speaking, a very general point in $Z(x)$ satisfies the same algebraic properties as $x$. In this paper, we study lengths of periodic points in whose definition (see §1.1) we need the norm map $|\cdot|:\mathbb{C}\to\mathbb{R}_{\geq 0}$ which is not algebraic. However $|\cdot|^{2}:\mathbb{C}\to\mathbb{R}$ is real algebraic. For this reason, we need to generalize the notion of transcendental points to respect the structure of $X_{K}(K)$ as an algebraic variety over $L$ and the structure of $X(\mathbb{C})$ as a real algebraic variety, and the generalized notion is important in the proof of Theorem 2.1.
The Weil restriction $R(X)$ of $X$ w.r.t. $\mathbb{C}/\mathbb{R}$ is an algebraic variety over $\mathbb{R}$ and the Weil restriction $R(X_{K})$ of $X_{K}$ w.r.t. $K/L$ is an algebraic variety over $L$. We have $R(X)=R(X_{K})\otimes_{L}\mathbb{R}$. Denote by $\pi_{L}:R(X)\to R(X_{K})$ the natural projection. For every $x\in X(\mathbb{C})$, let $Y(x)_{L}$ be the Zariski closure of $\pi_{L}(\psi_{\mathbb{C}}(x))$ in $R(X_{K})$ and $Y(x):=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, where $\psi_{\mathbb{C}}:X(\mathbb{C})\to R(X)(\mathbb{R})$ is the identification in §2.1. Set $Z^{\mathbb{R}}(x):=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$ which is a real Zariski closed subset of $X(\mathbb{C})$.
@@
$Z(x)\cap U=\{y\in U|\,\,h(y)=0\text{ for every }h\in\mathcal{O}^{K}(U)\text{ with }h(x)=0\}$
@@
As $\mathcal{O}(R(U)_{\mathbb{C}})=\mathcal{O}(R(U))\otimes_{\mathbb{R}}\mathbb{C}$, every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ can be viewed as a $\mathbb{C}$-valued algebraic function on $R(U)(\mathbb{R})$. Every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ induces a function $h\circ\psi_{\mathbb{C}}$ on $U(\mathbb{C})$. The functions of this form are exactly the $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Denote by $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$ the $\mathbb{R}$-algebra of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Since algebraic functions are real algebraic, we have a natural embedding $\mathcal{O}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$. By Proposition 2.8, we have
@@
Let $\mathcal{O}^{L}(R(U)):=\pi_{L}^{*}(\mathcal{O}(R(U_{K})))$ be the set of algebraic functions defined over $L$ on $R(U)$. Let $\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$ the image of $\mathcal{O}^{L}(R(U))\otimes_{L}K$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$, which is the set of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$ defined over $L$. It is clear that $\mathcal{O}^{K}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$. By Proposition 2.8, we have
@@
$Z^{\mathbb{R}}(x)\cap U(\mathbb{C})=\{y\in U(\mathbb{C})|\,\,h(y)=0\text{ for every }h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\text{ with }h(x)=0\}$
@@
Let $f_{K}:X^{\prime}_{K}\to X_{K}$ be a morphism between $K$-varieties. Set $X^{\prime}:=X^{\prime}_{K}\otimes_{K}\mathbb{C}$ and let $f:X^{\prime}\to X$ be the morphism induced by $f$. Let $x^{\prime}\in X^{\prime}(\mathbb{C})$ and $x:=f(x^{\prime})\in X(\mathbb{C})$. Then we have $f(Z^{\mathbb{R}}(x^{\prime}))\subseteq Z^{\mathbb{R}}(x)$.
@@
We have $Z^{\mathbb{R}}(x)\subseteq Z(x)$ and $Z^{\mathbb{R}}(x)$ is Zariski dense in $Z(x)$. In particular, if $x$ is transcendental, then $\dim_{\mathbb{R}}Z^{\mathbb{R}}(x)\geq 1$.
@@
$h\otimes_{\mathbb{C}}1\in\mathcal{O}(X)\otimes_{\mathbb{C}}\tau(\mathcal{O}(X))=\mathcal{C}^{\mathbb{R}-alg}(X)=\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}.$
View $\mathbb{C}$ as a vector space over $K$ and let $(e_{j})_{j\in J}$ be a $K$-basis of $\mathbb{C}$. We may assume that $0\in J$ and $1=e_{0}$. Write $h\otimes_{\mathbb{C}}1=\sum_{j\in J}g_{j}e_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}$, where $g_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)$. We have $g_{j}(x)=0$ for every $j\in J$.
@@
and $h\otimes_{\mathbb{C}}1$ is contained in the image of $\mathcal{O}(X)=\mathcal{O}^{K}(X)\otimes_{K}\mathbb{C}$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X).$ We have
@@
for every $n\neq 0$. So $g_{j}=\sum_{m\in N}b_{j,m,0}f_{m}\otimes 1\in\mathcal{O}^{K}(X)$. Since $g_{j}(x)=0$, $g_{j}|_{Z(x)}=0$. Then we have $h|_{Z(x)}=0$ which concludes the proof. ∎

### v3

@@ 1.6.1. The Zariski-dense orbit conjecture
By applying Corollary 1.7, we can give a new proof of a special case of the Zariski-dense orbit conjecture.
@@
Let $k$ be an algebraically closed field of characteristic 0. Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$. If $\{g\in k(X):g\circ f=g\}=k$, where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
@@
The converse of ZDO holds and is easy to prove. For some progress on ZDO, see for example [1, 2, 17, 27, 29].
@@
Let $X=\mathbb{P}^{1}_{k}\times\cdots\times\mathbb{P}^{1}_{k}$ be the product of $N$ copies of the projective line over an algebraically closed field $k$ of characteristic 0. Suppose that $f:X\to X$ is an endomorphism of the form $f_{1}\times\cdots\times f_{N}$, where $f_{j}:\mathbb{P}^{1}_{k}\to\mathbb{P}^{1}_{k}$ is a non-constant rational map for $1\leq j\leq N$. Then the ZDO holds for $(X,f)$.
@@
The original proof of Theorem 1.12 in [29] relies on the solution of the (adelic) Zariski-dense orbit conjecture for smooth projective surfaces [29, Theorem 1.15], the notion of adelic topology introduced in [29, §3], and a classification result [29, Proposition 9.2] on invariant subvarieties of $f:(\mathbb{P}^{1})^{N}\to(\mathbb{P}^{1})^{N}$ (see also [17, 8]). When $N=2$, Pakovich gave another proof [22] of Theorem 1.12 using his classification of invariant curves in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ and a height argument. We give a new proof in the case where all the $f_{i}$’s are non-exceptional of degree at least $2$; this proof does not require the ingredients mentioned above. Other cases of Theorem 1.12 are relatively easy or can be reduced to the case where all $f_{i}$ are non-exceptional, as shown in [29, §9.3].
@@ 1.6.2. A characterization of PCF maps
Here $K_{x}$ (resp. $L_{x}$) is any number field containing $\rho_{f}(x)$ (resp. $\lvert\rho_{f}(x)\rvert$), and $N_{K_{x}/\mathbb{Q}}$ (resp. $N_{L_{x}/\mathbb{Q}}$) is the norm map for the extension $K_{x}/\mathbb{Q}$ (resp. $L_{x}/\mathbb{Q}$), i.e., the determinant of the $\mathbb{Q}$-linear transformation induced by multiplication by $\rho_{f}(x)$ (resp. $|\rho_{f}(x)|$).
Clearly, the subspaces $V$ and $W$ above are independent of the choices of the fields $K_{x}$ and $L_{x}$, respectively. We also remark that the $\mathbb{Q}$-vector spaces $V$ in (2) and $W$ in (3) are the same.
@@
Similarly, the actual proof shows that we can weaken conditions (2) and (3) in Theorem 1.14, in the spirit of Remarks 1.3 and 1.6. Precisely, we can add two more equivalent statements to Theorem 1.14:
$(2)^{*}$ $\rho_{f}(x)\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $V^{*}=V^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $V^{*}$ is generated over $\mathbb{Q}$ by
@@
$(3)^{*}$ $\lvert\rho_{f}(x)\rvert\in\overline{\mathbb{Q}}$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$ and the $\mathbb{Q}$-subspace $W^{*}=W^{*}(f)$ of $\mathbb{R}$ is of finite dimension, where $W^{*}$ is generated over $\mathbb{Q}$ by
@@ 2. Rational maps with algebraic lengths
Let $K$ be an algebraically closed subfield of $\mathbb{C}$ which is invariant under complex conjugation $\tau$, i.e., $\tau(K)=K$. The aim of this section is to prove the following theorem.
@@
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and that for every $x\in{\rm Per}(f)(\mathbb{C})$, $\lvert\rho_{f}(x)\rvert\in K$. Then $f$ is defined over $K$.
@@
The condition in Theorem 2.1 can be weakened: it suffices to require that $f$ is not flexible Lattès and $\lvert\rho_{f}(x)\rvert\in K$ for all but finitely many $x\in{\rm Per}(f)(\mathbb{C})$. To show this, we need to replace $A$ in the proof in §2.7 by $A=(A_{n}=(RL^{*}(f)_{n})\cap K)_{n}$, and then the original proof applies. Here, $A_{n}=(RL^{*}(f)_{n})\cap K$ is the sub-multiset of the multiset $RL^{*}(f)_{n}$ consisting of elements that also lie in $K$; see §2.6 and §2.7.
Applying Theorem 2.1 to the case $K=\overline{\mathbb{Q}}$, we get Theorem 1.2. We introduce the notions of Weil restriction, admissible subsets, and transcendental points in §2.12.3. Then we construct certain moduli spaces and study the length map on them in §2.42.6. After these preparations, we complete the proof of Theorem 2.1 in §2.7.
@@ 2.1. Weil restriction
and both $\frac{u+\tau(u)}{2}$ and $\frac{u-\tau(u)}{2i}$ are contained in $L$. This completes the proof. ∎
To prove Theorem 2.1, we need to analyze the length map while respecting the structure of $\mathcal{M}_{d}(K)$ as an algebraic variety over $L$ using Weil restriction, due to the appearance of the absolute value in the condition. Hence, the notion of Weil restriction, which we will briefly recall, is necessary. See [23, §4.6] and [4, §7.6] for more details.
Denote by ${\rm Var}_{/K}$ (resp. ${\rm Var}_{/L}$) the category of varieties over $K$ (resp. $L$). For every variety $X$ over $K$, there is a unique variety $R(X)$ over $L$ which represents the functor ${\rm Var}_{/L}\to{\rm Sets}$ sending $V\in{\rm Var}_{/L}$ to ${\rm Hom}(V\otimes_{L}K,X)$. It is called the Weil restriction of $X$ with respect to $K/L$. The functor $X\mapsto R(X)$ is called the Weil restriction. There is a canonical bijection $\psi_{K}:X(K)\to R(X)(L)$ between sets. When $K=\mathbb{C}$, this map is a real analytic diffeomorphism. One may view $X(K)$ as the set of $L$-points of an $L$-algebraic variety $R(X)$ via $\psi_{K}$.
@@
The $L$-Zariski topology on $X(K)$ is the restriction of the Zariski topology on $R(X)$ to $R(X)(L)$ via the identification $\psi_{K}$. A subset $Y$ of $X(K)$ is called $L$-algebraic if it is closed in the $L$-Zariski topology. When $K=\mathbb{C}$, the $L$-Zariski topology coincides with the real Zariski topology as in [12, §8.1.1].
By the last statement of Proposition 2.6 below, the $L$-Zariski topology is finer than the Zariski topology on $X(K)$.
@@
Let $X,Y\in{\rm Var}_{/K}$. The following properties hold:
@@
If $X$ is the subvariety of $\mathbb{A}^{N}_{K}={\rm Spec}K[z_{1},\dots,z_{N}]$ defined by the equations $\sum_{I}a_{i,I}z^{I}=0$ ($i=1,\dots,s$), then $X^{\tau}$ is the subvariety of $\mathbb{A}^{N}_{K}$ defined by $\sum_{I}\tau(a_{i,I})z^{I}=0$ ($i=1,\dots,s$). The map $\tau:X=(X^{\tau})^{\tau}\to X^{\tau}$ sends a point $(z_{1},\dots,z_{N})\in X(K)$ to $(\tau(z_{1}),\dots,\tau(z_{N}))\in X^{\tau}(K)$.
@@
Proposition 2.8 ([23, Exercise 4.7]).
There is a canonical isomorphism
@@ 2.2. Admissible subsets
In this subsection, we recall the notion of admissible subsets of real algebraic varieties introduced in [12], which satisfies the important descending chain condition.
@@
Definition 2.9 ([12, §8.2]).
A closed subset $V$ of $X(\mathbb{R})$ (with the real topology) is called admissible if there is a morphism $f:Y\to X$ of real algebraic varieties and a Zariski closed subset $V^{\prime}\subseteq Y$ such that $V=f(V^{\prime}(\mathbb{R}))$ and $f$ is étale at every point in $V^{\prime}(\mathbb{R})$.
@@
Denote by $J$ the non-étale locus of $f$ in $Y$. Then $J\cap V^{\prime}(\mathbb{R})=\emptyset$. Since we may replace $(Y,V^{\prime})$ by $(Y\setminus J,V^{\prime}\cap(Y\setminus J))$, in the above definition we may further assume that $f$ is étale.
Proposition 2.11 ([12, Remarks 8.14, 8.15, and Proposition 8.16]).
The following basic properties hold:
@@
Theorem 2.12 ([12, Theorem 8.17]).
Let $(V_{n})_{n\geq 1}$ be a decreasing sequence of admissible subsets of $X(\mathbb{R})$. Then there exists $N\geq 1$ such that $V_{n}=V_{N}$ for all $n\geq N$.
@@ 2.3. Transcendental points
Let $X_{K}$ be a variety over $K$ and $X:=X_{K}\otimes_{K}\mathbb{C}$. We regard $X_{K}$ as a model of $X$ over $K$.
Denote by $\pi_{K}:X\to X_{K}$ the natural projection. For every point $x\in X(\mathbb{C})$, define $Z(x)_{K}$ to be the Zariski closure of $\pi_{K}(x)$ in $X_{K}$ and $Z(x):=\pi_{K}^{-1}(Z(x)_{K})\subseteq X$. It is clear that $Z(x)$ is irreducible. We call $Z(x)$ the $\mathbb{C}/K$-closure of $x$ with respect to the model $X_{K}$. We say that $x$ is transcendental if $\dim Z(x)\geq 1$ and call $\dim Z(x)$ the transcendence degree of $x$.
The notion of transcendental points plays an important role in the proof of some cases of the geometric Bombieri-Lang conjecture by Xie-Yuan [30, 31] and in the proof of the dynamical André-Oort conjecture for curves by Ji-Xie [11]. Roughly speaking, a very general point in $Z(x)$ satisfies the same algebraic properties as $x$. In this paper, we study lengths of periodic points, whose definition (see §1.1) involves the absolute value function $|\cdot|:\mathbb{C}\to\mathbb{R}_{\geq 0}$ which is not algebraic. However, the square $|\cdot|^{2}:\mathbb{C}\to\mathbb{R}$ is real algebraic. For this reason, we need to generalize the notion of transcendental points to respect the structure of $X_{K}(K)$ as an algebraic variety over $L$ and the structure of $X(\mathbb{C})$ as a real algebraic variety. This generalized notion is important in the proof of Theorem 2.1.
The Weil restriction $R(X)$ of $X$ with respect to $\mathbb{C}/\mathbb{R}$ is an algebraic variety over $\mathbb{R}$, and the Weil restriction $R(X_{K})$ of $X_{K}$ with respect to $K/L$ is an algebraic variety over $L$. We have $R(X)=R(X_{K})\otimes_{L}\mathbb{R}$. Denote by $\pi_{L}:R(X)\to R(X_{K})$ the natural projection. For every $x\in X(\mathbb{C})$, let $Y(x)_{L}$ be the Zariski closure of $\pi_{L}(\psi_{\mathbb{C}}(x))$ in $R(X_{K})$ and $Y(x):=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, where $\psi_{\mathbb{C}}:X(\mathbb{C})\to R(X)(\mathbb{R})$ is the identification from §2.1. Set $Z^{\mathbb{R}}(x):=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$, which is a real Zariski closed subset of $X(\mathbb{C})$.
@@
$Z(x)\cap U=\{y\in U\mid h(y)=0\text{ for every }h\in\mathcal{O}^{K}(U)\text{ with }h(x)=0\}$
@@
As $\mathcal{O}(R(U)_{\mathbb{C}})=\mathcal{O}(R(U))\otimes_{\mathbb{R}}\mathbb{C}$, every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ can be viewed as a $\mathbb{C}$-valued algebraic function on $R(U)(\mathbb{R})$. Every $h\in\mathcal{O}(R(U)_{\mathbb{C}})$ induces a function $h\circ\psi_{\mathbb{C}}$ on $U(\mathbb{C})$. Functions of this form are exactly the $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Denote by $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$ the $\mathbb{R}$-algebra of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$. Since algebraic functions are real algebraic, we have a natural embedding $\mathcal{O}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$. By Proposition 2.8, we have
@@
Let $\mathcal{O}^{L}(R(U)):=\pi_{L}^{*}(\mathcal{O}(R(U_{K})))$ be the set of algebraic functions defined over $L$ on $R(U)$. Let $\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$ be the image of $\mathcal{O}^{L}(R(U))\otimes_{L}K$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(U)$, which is the set of $\mathbb{C}$-valued real algebraic functions on $U(\mathbb{C})$ defined over $L$. It is clear that $\mathcal{O}^{K}(U)\subseteq\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)$. By Proposition 2.8, we have
@@
$Z^{\mathbb{R}}(x)\cap U(\mathbb{C})=\{y\in U(\mathbb{C})\mid h(y)=0\text{ for every }h\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(U)\text{ with }h(x)=0\}$
@@
Let $f_{K}:X^{\prime}_{K}\to X_{K}$ be a morphism between $K$-varieties. Set $X^{\prime}:=X^{\prime}_{K}\otimes_{K}\mathbb{C}$ and let $f:X^{\prime}\to X$ be the morphism induced by $f$. Let $x^{\prime}\in X^{\prime}(\mathbb{C})$ and $x:=f(x^{\prime})\in X(\mathbb{C})$. Then $f(Z^{\mathbb{R}}(x^{\prime}))\subseteq Z^{\mathbb{R}}(x)$.
@@
We have $Z^{\mathbb{R}}(x)\subseteq Z(x)$ and $Z^{\mathbb{R}}(x)$ is Zariski-dense in $Z(x)$. In particular, if $x$ is transcendental, then $\dim_{\mathbb{R}}Z^{\mathbb{R}}(x)\geq 1$.
@@
$h\otimes_{\mathbb{C}}1\in\mathcal{O}(X)\otimes_{\mathbb{C}}\tau(\mathcal{O}(X))=\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)=\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}.$
View $\mathbb{C}$ as a vector space over $K$ and let $(e_{j})_{j\in J}$ be a $K$-basis of $\mathbb{C}$. We may assume that $0\in J$ and $1=e_{0}$. Write $h\otimes_{\mathbb{C}}1=\sum_{j\in J}g_{j}e_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)\otimes_{K}\mathbb{C}$, where $g_{j}\in\mathcal{C}^{\mathbb{R}-{\rm alg},L}(X)$. Then $g_{j}(x)=0$ for every $j\in J$.
@@
and $h\otimes_{\mathbb{C}}1$ is contained in the image of $\mathcal{O}(X)=\mathcal{O}^{K}(X)\otimes_{K}\mathbb{C}$ in $\mathcal{C}^{\mathbb{R}-{\rm alg}}(X)$, we have
@@
for every $n\neq 0$. So $g_{j}=\sum_{m\in N}b_{j,m,0}f_{m}\otimes 1\in\mathcal{O}^{K}(X)$. Since $g_{j}(x)=0$, $g_{j}\mid_{Z(x)}=0$. Hence $h\mid_{Z(x)}=0$, which completes the proof. ∎